# Revisiting PSF models: unifying framework and high-performance implementation

### v1

@@ 5.1 PSF Gallery
We provide a gallery of PSFs to showcase different outputs of our library in Figure 3 and Figure 4. We use a wavelength of 632 nm, circular polarization for the vectorial models, and display the amplitude (in arbitrary units) of the beams for better contrast. Different 2D slices of the same 3D beam share the same dynamic range. We first compare the 3D unaberrated beams computed by the scalar and vectorial propagators for low- and high-NA. Two representative slices, one $x-y$ plane at the focus, the other $z-y$ plane through the center of the PSF, and vertical line profiles through the FOV center are shown in Figure 3. In the low-NA case, the resulting PSFs from both propagators are similar. The beams differ more from each other, as expected, in the high-NA case: the rings are blurred out as the energy is spread into different components of the focus field.
@@ 5.2.1 Speed Benchmark
We benchmark the runtime to compute a single $201\times 201$ 2D PSF image against the size of the pupil. The image is captured at focus and all input parameters take default values in our library. We compare the runtime of all four propagators on CPU and GPU and show the results in Fig. 5. The benchmarking was performed on a machine with an Intel i9-10900X CPU and an NVIDIA GeForce RTX 3090 GPU.
Figure 5: Speed benchmark. Time to generate a $201\times 201$ PSF for different numerical sizes of the pupil, for the different four propagators, on CPU (a) and GPU (b). Each data point in the plots is averaged over 10 runs.
We observe that the runtime of all propagators increases with the size of the pupil on CPU. The Cartesian propagators (red) are faster than the spherical ones (blue) on smaller sizes (<512 pixels) but slower on large sizes. This illustrates the difference in computational complexity of the different methods. Scalar propagators (dotted) are faster than their vectorial counterparts (solid), by roughly 1.5 times for Cartesian and 3 times for spherical cases. On GPU, Cartesian propagators are faster than on CPU at the same grid size and the curves (red) behave similarly; spherical propagators, however, exhibit a flat curve (blue) which indicates that they efficient benefit from the GPU parallelization. The speed improvement between scalar and vectorial propagators is small, especially for the Cartesian case. Hence, vectorial propagators should be preferred over the scalar ones as the accuracy gain does not come at a high computational cost. Moreover, Cartesian propagators are recommended if one works with images of small sizes for both scalar and vectorial cases; while spherical propagators should be preferred for large sizes, especially when GPU is available.
@@ 5.2.2 Accuracy Benchmark
We benchmark the accuracy of both Cartesian and spherical scalar propagators with the analytic Airy disk for asymptotic limit (the Fourier transform of a perfect circular aperture):
@@
The spherical integral is computed using two numerical integration rules: the Riemann rule (a first-order method) and the Simpson’s rule (a fourth-order method). The error is the $L_{2}$-norm of the difference between the output electric field $E$ of the propagator and $F_{\text{AD}}$
@@
Figure 6: Accuracy benchmark of the Cartesian propagator (red) and the spherical propagator using the Riemann rule (green) and Simpsons’ rule (blue). $h$ is the step size of integration.
We observe in Figure 6 that the error decreases as the number of points in the integration domain increases in all cases. The spherical propagator has a linear convergence rate using the Riemann rule (green) and 4th-order convergence rate using Simpsons’ rule (blue), which correlates well with their expected accuracy. The Cartesian propagator, which is based on our custom FFT, shows a convergence rate between first- and second-order.
@@ 6 Conclusion
In practice, the choice of propagators depends on the symmetry of the pupil function. Based on our benchmark, for axisymmetric pupil functions, spherical propagators are recommended thanks to their high accuracy and scalability. They are also particularly amenable to GPU parallelization. For non-axisymmetric cases, such as those involving Zernike aberrations or specialized phase masks, Cartesian propagators are required. In all these cases, the difference in computational time is relatively modest for pupil sizes up to a few hundreds. Hence, the vectorial Cartesian propagator could be a solid default choice in most applications.
@@
Vasiliki Stergiopoulou acknowledges funding from the Swiss National Science Foundation under Grant CRSII5_213521, "DigiLight - Programmable Third-Harmonic Generation (THG) Microscopy Applied to Advanced Manufacturing". Jonathan Dong acknowledges funding from the Swiss National Science Foundation (Grant PZ00P2_216211).

### v2

@@ 5.1 PSF Gallery
We provide in Figures 4 and 5 a gallery of PSFs to showcase our PyTorch implementation. We used a wavelength of 632 nm, a circular polarization for the vectorial models, and we display the amplitude (in arbitrary units) of the beams for better contrast. The 2D slices of the same 3D beam share the same dynamic range. In Figure 4, we present the focal electric field computed using both the scalar and vectorial models for objectives with different NAs. In the low-NA case (top two rows), the resulting PSFs from both propagators are similar. The beams differ more from each other, as expected, in the high-NA case (bottom two rows): the rings are blurred out as the energy is spread into different components of the focus field.
@@ 5.2.1 Speed Benchmark
We benchmark the runtime to compute a single $201\times 201$ 2D PSF image against the size of the pupil. The image is captured at focus and all input parameters take default values in our library. We compare the runtime of all four propagators on CPU and GPU and show the results in Figure 6. The benchmarking was performed on a machine with an Intel i9-10900X CPU and an NVIDIA GeForce RTX 3090 GPU.
Figure 6: Time to generate a $201\times 201$ PSF in terms of the numerical size of the pupil and the propagator, on CPU (a) and GPU (b). Each data point in the plots is averaged over 10 runs.
We observe that the runtime of all propagators increases with the size of the pupil on CPU. The Cartesian propagators (red) are faster than the spherical ones (blue) on small sizes (<512 pixels) but slower on larger sizes. This illustrates the computational complexity of each method. Scalar propagators (dotted) are faster than their vectorial counterparts (solid), by roughly 1.5 times for the Cartesian and 3 times for the spherical cases. On GPU, Cartesian propagators are faster than on CPU at the same grid size and the curves (red) behave similarly. Spherical propagators, however, exhibit a flat curve (blue), which indicates that they benefit heavily from the GPU parallelization. The speed improvement between scalar and vectorial propagators is small, especially for the Cartesian case. Hence, vectorial propagators should be preferred over the scalar ones as the accuracy gain does not come at a high computational cost. Moreover, Cartesian propagators are recommended if one works with images of small size for both scalar and vectorial cases. For larger sizes, however, spherical propagators should be preferred, especially when a GPU is available.
@@ 5.2.2 Accuracy Benchmark
We benchmark the accuracy of the Cartesian and spherical scalar propagators with the analytic Airy disk for asymptotic limit (the Fourier transform of a perfect circular aperture) given by
@@
The spherical integral is computed using two numerical integration rules: Riemann rule (a first-order method) and Simpson rule (a fourth-order method). The error $\delta$ is the $L_{2}$-norm of the difference between the output electric field $E$ of the propagator and $F_{\text{AD}}$, as in
@@
Figure 7: Accuracy of the Cartesian (red) and the spherical propagators using Riemann rule (green) and Simpson rule (blue). The step size of integration is $h$.
We observe in Figure 7 that the error decreases as the number of points in the integration domain increases in all cases. The spherical propagator has a linear convergence rate using Riemann rule (green) and 4th-order convergence rate using Simpson rule (blue), which correlates well with their expected accuracy. The Cartesian propagator, based on our custom FFT, shows a convergence rate between first- and second-order.
@@ 6 Conclusion
In practice, the choice of propagators depends on the symmetry of the pupil function. Based on our benchmark, for axisymmetric pupil functions, spherical propagators are recommended thanks to their high accuracy and scalability. They are also particularly amenable to GPU parallelization. For non-axisymmetric cases, such as those involving Zernike aberrations or specialized phase masks, Cartesian propagators are required. In all these cases, the difference in computational time is relatively modest for pupil sizes up to a few hundred pixels. Hence, the vectorial Cartesian propagator could be a solid default choice in most applications.
@@
Vasiliki Stergiopoulou acknowledges funding from the Swiss National Science Foundation under Grant CRSII5_213521, “DigiLight - Programmable Third-Harmonic Generation (THG) Microscopy Applied to Advanced Manufacturing.” Jonathan Dong acknowledges funding from the Swiss National Science Foundation (Grant PZ00P2_216211).
Acknowledgment
We would like to thank Eric Sinner for his valuable help in software engineering.